OpenStudy (ria23):

Can someone help me simplify the negative square root of -169?

OpenStudy (ria23):

\[-\sqrt{-169}\]

OpenStudy (sleepyjess):

have you learned about \(\normalsize\it i\) yet?

OpenStudy (studygurl14):

\[\sqrt{-1}=i\] So.... \[-\sqrt{-169} \rightarrow -i \sqrt{169}\]

OpenStudy (ria23):

@sleepyjess yea I have, but the negative outside of the square root was confusing me @StudyGurl14 I would just factor 169 from there right?

OpenStudy (sleepyjess):

@StudyGurl14 I was thinking it would be \(-\sqrt{169i}\)

OpenStudy (studygurl14):

No @sleepyjess because... \[-\sqrt{-169} \rightarrow -\sqrt{-1\times169} \rightarrow -\sqrt{-1}\sqrt{169} \rightarrow -i \sqrt{169}\]

OpenStudy (ria23):

@StudyGurl14 If I was to factor this out, would I end up with \[-13i \sqrt{13}\]

OpenStudy (ria23):

@sleepyjess I like yhur poptart owl.

OpenStudy (studygurl14):

Nope. since \[\sqrt{169} = 13\]

OpenStudy (studygurl14):

so... \[-i \sqrt{169} = -13i\]

OpenStudy (ria23):

Oh! >u< Thank yhu~

OpenStudy (studygurl14):

You're welcome